Find the distance between the parallel planes -2x - 2y+z=1 and -2.-2y+z=4 Decimal answers will not be accepted.
Answer #1

two parallel plomes ane. Pi : -2x - 24+2=1. = -22-24 +2-1 = 0. Pz: -2% -29+2 = 4. : -22-29+2-4=0. we know that distomced betso distance between given two parallel planes is d = 1